# Boosting Semi-Supervised 2D Human Pose Estimation by Revisiting Data Augmentation and Consistency Training

### v2

@@ Abstract
The 2D human pose estimation (HPE) is a basic visual problem. However, its supervised learning requires massive keypoint labels, which is labor-intensive to collect. Thus, we aim at boosting a pose estimator by excavating extra unlabeled data with semi-supervised learning (SSL). Most previous SSHPE methods are consistency-based and strive to maintain consistent outputs for differently augmented inputs. Under this genre, we find that SSHPE can be boosted from two cores: advanced data augmentations and concise consistency training ways. Specifically, for the first core, we discover the synergistic effects of existing augmentations, and reveal novel paradigms for conveniently producing new superior HPE-oriented augmentations which can more effectively add noise on unlabeled samples. We can therefore establish paired easy-hard augmentations with larger difficulty gaps. For the second core, we propose to repeatedly augment unlabeled images with diverse hard augmentations, and generate multi-path predictions sequentially for optimizing multi-losses in a single network. This simple and compact design is interpretable, and easily benefits from newly found augmentations. Comparing to SOTA approaches, our method brings substantial improvements on public datasets. Code is in https://github.com/hnuzhy/MultiAugs
Keywords:
@@
Semi-supervised learning Human pose estimation Data augmentation Consistency training
1 Introduction
@@
2D human pose estimation (HPE) aims to detect and represent human parts as sparse 2D keypoint locations in RGB images. It is the basis of many visual tasks such as action recognition [61, 13], person re-identification [71, 46] and 3D human shape regression [42, 41]. Modern data-driven HPE has been substantially improved by generous deep supervised learning approaches [4, 55, 49, 9, 59, 63, 35]. This greatly benefits from the collection and annotation of many large-scale public HPE datasets [1, 34, 53, 29]. However, compared to image classification and detection tasks requiring plain labels, obtaining accurate 2D keypoints from massive images is laborious and time-consuming. To this end, some researches [58, 37, 51, 23] try to alleviate this problem by introducing the semi-supervised 2D human pose estimation (SSHPE), which can subtly leverage extensive easier obtainable yet unlabeled 2D human images in addition to partial labeled data to improve performance. Although methods [58, 23] have improved the accuracy of SSHPE task, they overlooked two fundamental questions:
(a) Single Network
(b) Dual Networks
(c) Triple Networks
Figure 1: Frameworks of existing SSHPE methods including Single-Network (a) and Dual-Network (b) proposed by [58], Triple-Network (c) proposed by [23].
@@
Q1: How to judge the discrepancy of unsupervised data augmentations with different difficulty levels? In [58], as shown in Fig. 0(a), for a batch of unlabeled images $\mathbf{I}$, its easy augmentation $\mathbf{I}_{e}$ and hard augmentation $\mathbf{I}_{h}$ are generated separately. Then, predicted heatmap $\mathbf{H}_{e}$ of $\mathbf{I}_{e}$ is used as a pseudo label to teach the network to learn the harder counterpart $\mathbf{I}_{h}$ with its yielded heatmap $\mathbf{H}_{h}$. [58] finds that the large gap between two augmentations $(\mathbf{I}_{e},\mathbf{I}_{h})$ matters. Essentially, this is a pursuit for more advanced data augmentations in SSHPE. To rank augmentations of different difficulty levels, [58] observes precision degradation of a pretrained model by testing it on a dataset after corresponding augmentation. However, we declare that this manner is not rigorous. An obvious counter-case is that over-augmented samples approaching noise will get worst evaluation results, but such hard augmentations are meaningless. Contrastingly, we deem that a persuasive ranking requires independent training for each augmentation. We answer this question in detail in Sec. 3.1.
@@
Q2: How to generate multiple unsupervised signals for consistency training efficiently and concisely? Previous work [58] proposes to use a Single-Network to perform the unsupervised consistency training on the easy-hard pair $(\mathbf{I}_{e},\mathbf{I}_{h})$. It also gives a more complicated Dual-Network as in Fig. 0(b) for cross-training of two easy-hard pairs. SSPCM [23] even constructs a Triple-Network as in Fig. 0(c) for interactive training of three easy-hard pairs. This pattern of adding more networks with the increase of unsupervised signal pairs can certainly bring gains. But this is cumbersome and will decelerate the training speed proportionally. In practice, considering that augmentations are always performed on the same input, we can repeatedly augment $\mathbf{I}$ multiple times with $n$ diverse hard augmentations, and generate multi-path easy-hard pairs $\{(\mathbf{I}_{e},\mathbf{I}_{h_{1}}),(\mathbf{I}_{e},\mathbf{I}_{h_{2}}),...,(\mathbf{I}_{e},\mathbf{I}_{h_{n}})\}$. In this way, we can use only one single network (refer Fig. 6(a)) to optimize $n$ pairs of losses. This is also applicable to dual networks (refer Fig. 6(b)). We discuss this question in detail in Sec. 3.2.
@@
In this paper, we mainly revisit Q1 and Q2 to boost SSHPE. For Q1, after properly ranking existing basic augmentations [12, 70, 66, 10, 11], we naturally try to extend it to discover new strong augmentations through reasonable sequential combinations inspired by the AutoAug families [10, 33, 18, 72]. Rather than trivial enumeration, we notice the synergistic effects between different augmentations, and reveal novel paradigms for easily generating superior combinations. Our paradigms for the SSHPE task contain three principles: (P1) Do not combine MixUp-related augmentations. (P2) Try to utilize the synergistic effects. (P3) Do not over-combine too many augmentations. These principles have favorable interpretability, and bypass painstaking designs in other advanced augmentations [21, 38, 72, 43, 36, 17]. For Q2, we quantitatively validated the superiority of multi-path design over commonly used heatmaps fusion [45] and confidence masking [56, 23]. Combining it with newly found advanced augmentations, our Single-Network based approach can surpass the original Dual-Network [58] and come close to SSPCM using a Triple-Network. Our contributions are as follows:
@@
  • We comprehensively evaluated the difficulty levels of existing data augmentations suitable for the SSHPE task, validated their synergistic effects by properly combining different basic augmentations, and presented novel combination paradigms which are intuitively interpretable.
@@
  • We proposed to generate multi-path predictions of separately strongly augmented samples for training only one single model, rather than adding auxiliary networks. Thus, we can optimize multiple unsupervised losses efficiently and concisely, and benefit from distinct superior augmentations.
@@
  • We achieved new SOTA results on public SSHPE benchmarks with less training time and parameters under same settings of previous methods.
2 Related Work
@@
Semi-Supervised Learning (SSL). SSL is invented for improving the classification task by exploiting a small set of labeled data and a large set of unlabeled data. It can be categorized into pseudo-label (PL) based [45, 40, 57, 47, 16, 52] and consistency-based [27, 50, 3, 56, 67, 14, 15]. PL-based methods iteratively add unlabeled images into the training data by pseudo-labeling them with a pretrained or gradually enhanced model. It needs to find suitable thresholds to mask out uncertain samples with low-confidence, which is a crucial yet tricky issue. Consistency-based methods enforce model outputs to be consistent when its input is randomly perturbed. They have shown to work well on many tasks. For example, MixMatch [3] combines the consistency regularization with the entropy minimization to obtain confident predictions. FixMatch [47] utilizes a weak-to-strong consistency regularization and integrates the pseudo-labeling to leverage unlabeled data. FlexMatch [67] and FreeMatch [52] adopt the curriculum learning and adaptive thresholding based on FixMatch, respectively. CRMatch [14] and SAA [15] try to design strategies and augmentations to enhance the consistency training. These SSL methods give us primitive inspirations.
@@
Semi-Supervised Human Pose Estimation (SSHPE). The SSHPE task is relatively less-studied comparing to other visual tasks such as classification and object detection. A few representative methods are based on pseudo labeling [51, 48] or consistency training [58, 37, 28, 23]. SSKL [37] designs a semantic keypoint consistency constraint to learn invariant representations of same keypoints. It has been evaluated on small-scale HPE benchmarks MPII [1] and LSP [24], instead of the larger COCO [34]. Following it, PLACL [51] introduces the curriculum learning by auto-selecting dynamic thresholds for producing pseudo-labels via reinforcement learning. Inspired by co-training [44] and dual-student [26], Dual-Network [58] points out the typical collapsing problem in SSHPE, and proposes the easy-hard augmentation pair on the same input to imitate teacher-student signals without relying on Mean Teacher [50]. SSPCM [23] extends the Dual into Triple by adding an auxiliary teacher for interactive training in multi-steps. It designs a handcrafted pseudo-label correction based on the predicted position inconsistency of two teachers, and has achieved SOTA performances. We revisit the less efficient-consistency training way in [58, 23], and propose to upgrade the Single-Network furtherly by multi-path predictions.
@@
Unsupervised Data Augmentations. The UDA [56] has emphasized and verified the key role of high-quality noise injection (e.g., data augmentations) in improving unsupervised consistency training. It utilizes advanced augmentations [10, 11] to promote the SSL classification. Then, [58] transfers the positive correlation between strong augmentations and SSL performance to the HPE field. It introduces a more advanced augmentation called Joint Cutout inspired by Cutout [12]. Similarly, SSPCM [23] provides a harder keypoints-sensitive augmentation Cut-Occlude inherited from CutMix [66]. In this paper, we thoroughly revisit existing data augmentations suitable to SSHPE, give a rank of their difficulty levels by controlled trainings, and produce simple paradigms for getting novel superior joint-related augmentations. We also compare them with other well-designed counterparts [11, 38, 17] to reveal our advantages.
3 Empirical Studies
@@
Problem Definition: The task of 2D HPE is to detect $k$ body joints in an image $\mathbf{I}\in\mathbb{R}^{h\times w\times 3}$. The state-of-the-art methods tend to estimate $k$ Gaussian heatmaps $\mathbf{H}\in\mathbb{R}^{\frac{h}{s}\times\frac{w}{s}\times k}$ downsampled $s$ times. For inference, each keypoint is located by finding the pixel with largest value in its predicted heatmap. We denote the labeled and unlabeled training sets as $\mathcal{D}^{l}\!=\!\{(\mathbf{I}^{l}_{i},\mathbf{H}^{l}_{i})\}|^{N}_{i=1}$ and $\mathcal{D}^{u}\!=\!\{\mathbf{I}^{u}_{i}\}|^{M}_{i=1}$, respectively. The $\mathbf{H}^{l}$ are ground-truth heatmaps generated using 2D keypoints. For supervised training of the network $f$, we calculate the MSE loss:
$\small\mathcal{L}_{s}=\mathbb{E}_{\mathbf{I}\in\mathcal{D}^{l}}||f(T_{e}(\mathbf{I}))-T_{e}(\mathbf{H})||^{2}~{}$ (1)
@@
$\small\mathcal{L}_{u}=\mathbb{E}_{\mathbf{I}\in\mathcal{D}^{u}}||T_{e\rightarrow h}(f(T_{e}(\mathbf{I})))-f(T_{h}(\mathbf{I}))||^{2}~{}$ (2)
@@
where $T_{h}$ is a harder augmentation with strong perturbations than affine-based $T_{e}$. The $T_{e\rightarrow h}$ means a known affine transformation on heatmaps if $T_{h}$ contains additional rotation and scaling operations. In this way, we can obtain a paired easy-hard augmentations $(\mathbf{I}_{e},\mathbf{I}_{h})\!=\!(T_{e}(\mathbf{I}),T_{h}(\mathbf{I}))$ for generating corresponding teacher signals and student signals. During training, we should stop gradients propagation of teacher signals to avoid collapsing.
Next, we answer two questions Q1 and Q2 by extensive empirical studies in Sec. 3.1 and Sec. 3.2, respectively. After that, we provide a perspective for understanding the pursuit of strong augmentations in Sec. 3.3.
3.1 Paradigms of Generating Superior Augmentations
@@
Figure 2: Comparison of applying different easy-hard pairs for training a Single-Network model as in Fig. 0(a). We can sort these six augmentations indisputably based on either best mAP results or distinct convergence curves.
Figure 2: Comparison of applying different easy-hard pairs for training a Single-Network model as in Fig. 0(a). We can sort these six augmentations indisputably based on either best mAP results or distinct convergence curves.
Figure 3: Illustrations of two novel superior combinations $T_{JOCO}$ and $T_{JCCM}$. Either of them is a sequential operations using ready-made collaborative augmentation. And $T_{JO}$ and $T_{CM}$ introduce extra patches cropped from other images which are not displayed.

### v3

@@ Abstract
The 2D human pose estimation (HPE) is a basic visual problem. However, its supervised learning requires massive keypoint labels, which is labor-intensive to collect. Thus, we aim at boosting a pose estimator by excavating extra unlabeled data with semi-supervised learning (SSL). Most previous SSHPE methods are consistency-based and strive to maintain consistent outputs for differently augmented inputs. Under this genre, we find that SSHPE can be boosted from two cores: advanced data augmentations and concise consistency training ways. Specifically, for the first core, we discover the synergistic effects of existing augmentations, and reveal novel paradigms for conveniently producing new superior HPE-oriented augmentations which can more effectively add noise on unlabeled samples. We can therefore establish paired easy-hard augmentations with larger difficulty gaps. For the second core, we propose to repeatedly augment unlabeled images with diverse hard augmentations, and generate multi-path predictions sequentially for optimizing multi-losses in a single network. This simple and compact design is interpretable, and easily benefits from newly found augmentations. Comparing to state-of-the-art SSL approaches, our method brings substantial improvements on public datasets. And we extensively validate the superiority and versatility of our approach on conventional human body images, overhead fisheye images, and human hand images. The code is released in https://github.com/hnuzhy/MultiAugs.
Index Terms:
@@
Semi-supervised learning, human pose estimation, data augmentation, consistency training
I Introduction
@@
The 2D human pose estimation (HPE) aims to detect and represent human parts as sparse 2D keypoint locations in RGB images. It is the basis of many visual tasks such as action recognition [1, 2], person re-identification [3, 4], 3D pose lifting [5, 6] and 3D human shape regression [7, 8]. Modern data-driven HPE has been substantially improved by generous deep supervised learning approaches [9, 10, 11, 12, 13]. This greatly benefits from the collection and annotation of many large-scale public HPE datasets [14, 15, 16]. However, compared to image classification and detection tasks requiring plain labels, obtaining accurate 2D keypoints from massive images is laborious and time-consuming. To this end, some researches [17, 18, 19, 20, 21] try to alleviate this problem by introducing the semi-supervised 2D human pose estimation (SSHPE), which can subtly leverage extensive easier obtainable yet unlabeled 2D human images in addition to partial labeled data to improve performance. Although existing methods [17, 20] have improved the accuracy of SSHPE task, they overlooked two fundamental questions:
@@
Q1: How to judge the discrepancy of unsupervised data augmentations with different difficulty levels? As shown in Fig. 1a, for a batch of unlabeled images $\mathbf{I}$, its easy augmentation $\mathbf{I}_{e}$ and hard augmentation $\mathbf{I}_{h}$ are generated separately in [17]. Then, predicted heatmap $\mathbf{H}_{e}$ of $\mathbf{I}_{e}$ is used as a pseudo label to teach the network to learn the harder counterpart $\mathbf{I}_{h}$ with its yielded heatmap $\mathbf{H}_{h}$. [17] finds that the large gap between two augmentations $(\mathbf{I}_{e},\mathbf{I}_{h})$ matters. Essentially, this is a pursuit for more advanced data augmentations in SSHPE. To rank augmentations of different difficulty levels, [17] observes precision degradation of a pretrained model by testing it on a dataset after corresponding augmentation. However, we declare that this manner is not rigorous. An obvious counter-case is that over-augmented samples approaching noise will get worst evaluation results, but such hard augmentations are meaningless. Contrastingly, we deem that a persuasive ranking requires independent training for each augmentation. We answer this question in detail in Sec. III-A.
@@
Q2: How to generate multiple unsupervised signals for consistency training efficiently and concisely? Previous work [17] proposes to use a Single-Network to perform the unsupervised consistency training on the easy-hard pair $(\mathbf{I}_{e},\mathbf{I}_{h})$. It also gives a more complicated Dual-Network as in Fig. 1b for cross-training of two easy-hard pairs. SSPCM [20] even constructs a Triple-Network as in Fig. 1c for interactive training of three easy-hard pairs. This pattern of adding more networks with the increase of unsupervised signal pairs can certainly bring gains. But this is cumbersome and will decelerate the training speed proportionally. In practice, considering that augmentations are always performed on the same input, we can repeatedly augment $\mathbf{I}$ multiple times with $n$ diverse hard augmentations, and generate multi-path easy-hard pairs $\{(\mathbf{I}_{e},\mathbf{I}_{h_{1}}),(\mathbf{I}_{e},\mathbf{I}_{h_{2}}),...,(\mathbf{I}_{e},\mathbf{I}_{h_{n}})\}$. In this way, we can use only one single network (refer Fig. 7a) to optimize $n$ pairs of losses. This is also applicable to dual networks (refer Fig. 7b). We discuss this question in detail in Sec. III-B.
(a) Single-Network
(b) Dual-Network
(c) Triple-Network
Figure 1: Frameworks of existing semi-supervised human pose estimation (SSHPE) methods including (a) Single-Network and (b) Dual-Network which is originally proposed by [17], and (c) Triple-Network that is proposed by [20].
@@
In this paper, we mainly revisit Q1 and Q2 to boost SSHPE. For Q1, after properly ranking existing basic augmentations [22, 23, 24, 25, 26], we naturally try to extend it to discover new strong augmentations through reasonable sequential combinations inspired by the AutoAug families [25, 27, 28, 29]. Rather than trivial enumeration, we notice the synergistic effects between different augmentations, and reveal novel paradigms for easily generating superior combinations. Our paradigms for the SSHPE task contain three principles: (P1) Do not combine MixUp-related augmentations. (P2) Try to utilize the synergistic effects. (P3) Do not over-combine too many augmentations. These principles have favorable interpretability, and bypass painstaking designs in other advanced augmentations [30, 31, 29, 32, 33, 34]. For Q2, we quantitatively validated the superiority of multi-path design over commonly used heatmaps fusion [35] and confidence masking [36, 20]. Combining it with newly found advanced augmentations, our Single-Network based approach can surpass the original Dual-Network [17] and come close to SSPCM using a Triple-Network.
In summary, our contributions are three-folds:
@@
  • We comprehensively evaluated difficulty levels of existing data augmentations suitable for the SSHPE task, validated their synergistic effects by properly combining different basic augmentations, and presented novel combination paradigms which are intuitively interpretable.
@@
  • We proposed to easily generate multi-path predictions of strongly augmented samples in separate for training only one single model, rather than adding more auxiliary networks. Therefore, we can optimize multiple unsupervised losses efficiently and concisely, and benefit from distinct superior augmentations.
@@
  • We achieved new state-of-the-art results on public SSHPE benchmarks (ranging from conventional human body images to overhead fisheye images or conventional human hand images) with less training time and parameters under same settings of previous methods.
Our paper is organized as follows. In Sec. II, we introduce various related works, including semi-supervised learning (SSL), semi-supervised human pose estimation (SSHPE), and unsupervised data augmentation techniques. In Sec. III, we first give a formal definition of SSHPE. Then, we focus on answering the raised two core questions in the form of empirical studies, along with theoretical analysis of the effectiveness of advanced augmentations. In Sec. IV, we explain the overall training framework of the proposed MultiAugs using the Single-Network or Dual-Network, which has wide adaptability and versatility. Finally, in Sec. V, we present a large number of experimental results, including quantitative and qualitative comparisons of our MultiAugs with other methods on multiple HPE datasets, as well as sufficient ablation studies.
II Related Work
II-A Semi-Supervised Learning (SSL)
@@
SSL is originated in the classification task by exploiting a small set of labeled data and a large set of unlabeled data. It can be categorized into pseudo-label (PL) based [35, 37, 38, 39, 40, 41] and consistency-based [42, 43, 44, 36, 45, 46, 47]. PL-based methods iteratively add unlabeled images into the training data by pseudo-labeling them with a pretrained or gradually enhanced model. It needs to find suitable thresholds to mask out uncertain samples with low-confidence, which is a crucial yet tricky issue. Consistency-based methods enforce model outputs to be consistent when its input is randomly perturbed. They have shown to work well on many tasks. For example, MixMatch [44] combines the consistency regularization with the entropy minimization to obtain confident predictions. FixMatch [39] utilizes a weak-to-strong consistency regularization and integrates the pseudo-labeling to leverage unlabeled data. FlexMatch [45] and FreeMatch [41] adopt the curriculum learning and adaptive thresholding based on FixMatch, respectively. CRMatch [48] and SAA [46] try to design strategies and augmentations to enhance the consistency training. These SSL methods give us primitive inspirations.
II-B Semi-Supervised Human Pose Estimation (SSHPE)
@@
This is relatively less-studied comparing to other visual tasks like classification and object detection. A few SSHPE methods are based on pseudo labeling [19, 49, 21] or consistency training [17, 18, 50, 20, 51]. SSKL [18] designs a semantic keypoint consistency constraint to learn invariant representations of same keypoints. It is evaluated on small-scale HPE datasets MPII [14] and LSP [52], instead of the larger COCO [15]. Following it, PLACL [19] introduces the curriculum learning by auto-selecting dynamic thresholds for producing pseudo-labels via reinforcement learning. Inspired by co-training [53] and dual-student [54], Dual-Network [17] points out the typical collapsing problem, and proposes the easy-hard augmentation pair on the same input to imitate teacher-student signals without relying on Mean Teacher [43]. SSPCM [20] extends the Dual into Triple by adding an auxiliary teacher for interactive training in multi-steps. It designs a handcrafted pseudo-label correction based on the predicted position inconsistency of two teachers, and has achieved SOTA performances. Still based on Dual-Network, Pesudo-HMs [21] utilizes the cross-student uncertainty to propose a threshold-refine procedure, which can select reliable pseudo-heatmaps as targets for learning from unlabeled data. While, in this paper, we revisit the less efficient consistency training in [17, 20], and propose to upgrade the Single-Network by exploring multi-path predictions.
II-C Unsupervised Data Augmentations
@@
The UDA [36] has emphasized and verified the key role of high-quality noise injection (e.g., data augmentations) in improving unsupervised consistency training. It utilizes advanced augmentations [25, 26] to promote the SSL classification. Then, [17] transfers the positive correlation between strong augmentations and SSL performance to the HPE field. It introduces a more advanced augmentation called Joint Cutout inspired by Cutout [22]. Similarly, SSPCM [20] provides a harder keypoints-sensitive augmentation Cut-Occlude inherited from CutMix [24]. In this paper, we thoroughly revisit existing data augmentations suitable to SSHPE, give a rank of their difficulty levels by controlled training, and produce simple paradigms for getting novel superior joint-related augmentations. We also compare them with other well-designed counterparts [26, 31, 34] to reveal our advantages.
III Empirical Studies
@@
Problem Definition: The task of 2D HPE is to detect $k$ body joints in an image $\mathbf{I}\in\mathbb{R}^{h\times w\times 3}$. The state-of-the-art methods [55, 56] tend to estimate $k$ Gaussian heatmaps $\mathbf{H}\in\mathbb{R}^{\frac{h}{s}\times\frac{w}{s}\times k}$ downsampled $s$ times. For inference, each keypoint is located by finding the pixel with largest value in its predicted heatmap. We denote the labeled and unlabeled training sets as $\mathcal{D}^{l}\!=\!\{(\mathbf{I}^{l}_{i},\mathbf{H}^{l}_{i})\}|^{N}_{i=1}$ and $\mathcal{D}^{u}\!=\!\{\mathbf{I}^{u}_{i}\}|^{M}_{i=1}$, respectively. Here, the $\mathbf{I}^{l}_{i}$ or $\mathbf{I}^{u}_{i}$ means a labeled or unlabeled image sample, respectively. And $N$ or $M$ is the total number of image samples. The $\mathbf{H}^{l}$ are ground-truth heatmaps generated using 2D keypoints. For supervised training of the network $f$, we calculate the MSE loss:
$\mathcal{L}_{s}=\mathbb{E}_{\mathbf{I}\in\mathcal{D}^{l}}||f(T_{e}(\mathbf{I}))-T_{e}(\mathbf{H})||^{2},~{}$ (1)
@@
$\mathcal{L}_{u}=\mathbb{E}_{\mathbf{I}\in\mathcal{D}^{u}}||T_{e\rightarrow h}(f(T_{e}(\mathbf{I})))-f(T_{h}(\mathbf{I}))||^{2},~{}$ (2)
@@
where $T_{h}$ is a harder augmentation with strong perturbations than affine-based $T_{e}$. The $T_{e\rightarrow h}$ means a known affine transformation on heatmaps if $T_{h}$ contains additional rotation and scaling operations. In this way, we can obtain a paired easy-hard augmentations $(\mathbf{I}_{e},\mathbf{I}_{h})\!=\!(T_{e}(\mathbf{I}),T_{h}(\mathbf{I}))$ for generating corresponding teacher signals and student signals. During training, we should stop gradients propagation of teacher signals to avoid collapsing. Next, we answer two questions Q1 and Q2 by extensive empirical studies in Sec. III-A and Sec. III-B, respectively. After that, we provide a theoretical perspective of pursuiting stronger augmentations in Sec. III-C.
III-A Paradigms of Generating Superior Augmentations
@@
Figure 2: Comparison of applying different easy-hard pairs for training a Single-Network model as in Fig. 1a. We can rank these six augmentations indisputably based on either best mAP results or distinct convergence curves.
III-A1 Ranking of Basic Augmentations